# Variation in the size of the Photon Sphere and Black Hole Shadow in the Modified Gravity

### v2

@@ Abstract
Black hole shadows are a widespread topic in astrophysics. This paper searches for an optical view of the black hole and the relationship between black hole shadow and photon sphere with curvature. We were inspired by the observations of Sagittarius $A^{*}$ and supermassive black hole M87* through the Event Horizon Telescope. We have found the signature of the modified theory of gravity on the photon sphere and shadow. In this paper, we considered $f(R)$ modified theory of gravity. Our finding is that the new scalar degree of freedom $F$ appears in the expressions of photon sphere and Black hole shadow. Both are changed with $F$. and indeed, they depends on spacetime metric.
I introduction
Black holes are the mysterious and most complex structure in universe. Therefore, researcher want explore the this astrophysical object. There are many areas of the black hole. This article seeks to unravel the complexities surrounding black holes in the photon sphere snepppen2021divergent , the very first layer of these bodies. These are the boundaries of the Black Hole Shadow (BHS).
@@
The first image of the shadow was captured in 2019 by EHT EventHorizonTelescope:2019ggy . This is a groundbreaking result for studying the nature of black hole. The observational result of EHT is useful for testing the theory of gravity.
@@
Till now, we do not have direct observational evidences about the dark matter particle. Another mysterious component of the universe is dark energy, which responsible for the accelerated expansion of the universe. There are many methods and theories to solve both mysteries, dark matter and dark energy. The simplest way to tackle these problems by introducing a $\Lambda$ term for late time expansion Riess:1998fmf ; Carroll:2000fy ; Sahni:2002kh ; Sahni:1999gb ; Sahni:2015hbf and Cold Dark Matter (CDM) as a candidate of dark matter Bertone:2004pz ; Armendariz-Picon:2013jej ; Ilic:2020onu ; Perivolaropoulos:2021jda ; Yang:2025ume . There are many other way to explain them and one of them is the modification in the curvature. We used $f(R)$ models for accelerated expansion and large scale structure formation Sharma:2022fiw ; Sharma:2022tce and dark matter Yadav:2018llv ; Sharma:2019yix ; Sharma:2020vex .
@@
Shadow is like an illusion of four dimensions of the space-time and it depends on the many factors like background metric, static or moving observer and its position etc. Recent picture taken by the Event Horizon Telescope (EHT) of the Sagittarius $A^{*}$ (Sgr A*, pronounced ”sadge-ay-star”) EHT:2022 .
@@
The trajectory of light bends around the object because of high-density. If light bends so strongly that it moves on circular path because of a high density object like a black hole snepppen2021divergent . These orbits of light are known as photon spheres. Dark spots in the middle are termed as black hole shadow. In 1966, Synge studied path of the photons for null geodesics synge1966escape for Schwarzschild and non-rotating black hole.
@@
In past few years, many researchers have been investigated the black hole shadows. We know that the size of the black hole shadow is approximately 2.4 times larger than the particle horizon Wang:2022kvg ; Perlick2021CalculatingBH . This introduces some concept about the illusion of black hole shadow. It can be explained with the help of location of the observer. That means BHS depends on observer’s position, accretion disc that means matter around the horizon, rotation of black hole, and charged black hole takahashi2004shapes ; li2020shadow ; gralla2019black . Shapes of BHS strongly depends on modified theory of gravity models. Black hole shadow can be strong feature of the BH that constraint model’s parameters of the modified theory of gravity.
This paper organized as, in the first we introduced black hole shadow, photon sphere and observational values of the parameters. Second describe about the theoretical background of the photon sphere and black hole shadow. The generalized formulae established in the third section and also calculate photon sphere and black hole shadow for Schwarzschild, Kottler and Kerr spacetime metric. Fourth is based on the our work in modified theory of gravity and last section, we have discussed our outcomes.
I.1 Observational Evidence
The image of the M87 black hole taken by the Event Horizon Telescope (EHT) 2024A&A…681A..79E reveals that the angular radius EventHorizonTelescope:2019dse $42\pm 3$ micro arc-seconds ($\mu as$), and angular gravitational radius 3.8 $\pm$ 0.4 $\mu$ac, mass $M=(6.5\pm 0.7)\times 10^{9}$ M an asymmetric bright emission ring with a diameter (shadow angular diameter) of $42\pm 3$ $\mu as$, exhibiting a deviation from circularity $\Delta C\leq 0.1$ and an axis ratio $\lesssim 4/3$. This data is compatible with Kerr black hole in GR background.
@@ I.1 Observational Evidence
Another observational data released by the Event Horizon Telescope (EHT) akiyama2022first of the supermassive black hole Sgr A, showing an angular shadow diameter $\theta_{sh}=48.7\pm 0.7$ $\mu as$ and Schwarzschild shadow deviation $-0.08^{+0.09}_{-0.09}$ (VLTI), $-0.04^{+0.09}_{-0.10}$ (Keck) with mass of the black hole $M=4.0^{+1.1}_{-0.6}\times 10^{6}$ $M_{\odot}$ EventHorizonTelescope:2019ggy .
@@ II Theory background
The black hole shadow and photon sphere can be calculated analytically in general theory of relativity or other theories. These depend on the metric tensor and geodesic for photon. Now, we have the Schwarzschild metric for static, non-rotating, and uncharged symmetric sphere
@@
Non zero metric tensor components are
$g_{tt}$ $=-\Big{(}1-\frac{2GM}{c^{2}r}\Big{)},$ $g_{rr}$ $=\Big{(}1-\frac{2GM}{c^{2}r}\Big{)}^{-1}$
$g_{\theta\theta}$ $=r^{2},$ $g_{\phi\phi}$ $=r^{2}sin^{2}\theta.$
We have the Lagrangian
@@
Euler-Lagrange equation is
@@
Lagrangian have two cyclic coordinates $t$ and $\phi$ therefore we have two constant of motion ${\partial\mathcal{L}}/{\partial t}=0$ and ${\partial\mathcal{L}}/{\partial\phi}=0$. Therefore, we have two conserved quantities $E$ and $J_{z}$
@@
where $J_{z}$ is the angular momentum of the photon.
@@ II.1 Geodesic and photon orbit
Einstein’s general theory of relativity revolutionized our grasp of gravity, depicting it not as a force but as a curvature of spacetime. Imagine spacetime as a flexible fabric; massive objects like stars and planets create indentations, and black holes represent points where the fabric is stretched into an infinitesimal point-a singularity.
General theory of relativity is the most successful theory of universe. This theory explains many phenomena of the universe. We used Schwarzschild metric equation (1) to calculate photon sphere
Now we have calculated about the mathematical and geometrical expression of the photon sphere. Photons move around the massive object like black hole on the null geodesic i.e. $ds^{2}=0$ or $g_{\mu\nu}dx^{\mu}dx^{\nu}=0$. texbook
Null geodesic equation in the equatorial plane is
$-\Big{(}1-\frac{R_{s}}{r}\Big{)}dt^{2}+\frac{dr^{2}}{\Big{(}1-\frac{R_{s}}{r}\Big{)}}+r^{2}d\phi^{2}=0.$ (7)
If we divide by the affine parameter $d\lambda$
$-\Big{(}1-\frac{R_{s}}{r}\Big{)}\dot{t}^{2}+\frac{\dot{r}^{2}}{\Big{(}1-\frac{R_{s}}{r}\Big{)}}+r^{2}\dot{\phi}^{2}=0.$ (8)
@@
Figure 1: The dotted line is known as the angular diameter of black hole shadow $\alpha_{sh}$ or acceptance angle $r_{sh}=3\sqrt{3}m$ and red circle is the photon sphere $r_{ph}=3m$ for static observer. If observer is at the large distance $r_{o}>>m$, then $r_{sh}$ equal to impact factor Wang:2022kvg .
Figure 2: This figure shows that the sizes of black hole horizon, photon sphere and black hole shadow, respectively in Schwarzschild spacetime metric and for static observer Wang:2022kvg ; Perlick2021CalculatingBH
@@ II.2 Effective Potential
The effective potential can be used to calculate the radius of the photon sphere by $dV_{eff}/dr=0$. Now calculate the effective potential using the null geodesic equation (8) and using the value of $E$\&& $J_{z}$ from equations (5) and (6), respectively.
$-\frac{E^{2}}{\Big{(}1-\frac{R_{s}}{r}\Big{)}}+\frac{\dot{r}^{2}}{\Big{(}1-\frac{R_{s}}{r}\Big{)}}+\frac{J^{2}}{r^{2}}=0,$ (9)
@@ II.3 Photon Sphere
Radius of the photon sphere calculated from null geodesic, i.e., the geodesic for photon in the equatorial plane. The radius of the photon sphere calculated by using effective potential (11) and differentiating with respect to $r$
$\frac{dV_{eff}}{dr}=-\frac{J^{2}}{r^{3}}+\frac{3}{2}\frac{J^{2}}{r^{4}}.$ (12)
@@
$-\frac{J^{2}}{r^{3}}+\frac{3}{2}\frac{J^{2}}{r^{4}}=0,$ (13)
@@ II.4 Shadow of Black hole
Mathematical explanation of the black hole shadow for the first time given by Synge Synge:1966 in 1996. He considered spherical symmetric Schwarzschild metric foe static universe. He takes $\rho=r/2m$ and $\tau=t/2m$, where $G=c=1$. Constant of motion for angular momentum and energy conservation given as $\rho^{2}\dot{\theta}=\alpha$ and $(1-\rho^{-1})\dot{t}=\alpha\beta$, respectively. Path of the photon $\theta=$ const. &\&& ϕ=italic-$\phi=$const. and Small deviation of light ray (photon) is $\cot\alpha=\sqrt{g_{\rho\rho}/g_{\theta\theta}}(d\rho/d\theta)$ and $\sin^{2}\alpha=(\rho-1)/\rho^{3}\beta^{2}$. We have $\beta=4/27$ for $\rho=3/2$.
@@ III General expression of Black Hole Shadow and photon sphere
We have find out the general solution of the black hole shadow for spherical symmetric case in the static universe. In this case metric is
@@
where, $A(r)$, $B(r)$ and $D(r)$ are constant. If we use $A(r)=1-2m/r$, $B(r)=(1-2m/r)^{-1}$ and $D(r)=r^{2}$, $m=G_{N}M/c^{2}$.
We find constant of motion equation form (4) as
$-p_{t}=E=-A(r)\dot{t};$ $p_{\theta}=J_{z}=D(r)\sin^{2}\theta\dot{\phi},$ (19)
@@
$E=-A(r)\dot{t};$ $J_{z}=D(r)\dot{\phi}.$ (20)
We have effective potential $V_{eff}$ from Hamiltonian Wang:2022kvg , $\mathcal{H}=(1/2)g^{\mu\nu}p_{\mu}p_{\nu}$
$\mathcal{H}=\frac{1}{2}\Big{(}\frac{p_{t}^{2}}{A(r)}+\frac{p_{r}^{2}}{B(r)}+\frac{p_{\theta}^{2}}{D(r)}+\frac{p_{\phi}^{2}}{D(r)\sin^{2}\theta}\Big{)},$ (21)
$\mathcal{H}=\frac{1}{2}\Big{(}-\frac{E^{2}}{A(r)}+\frac{p_{r}^{2}}{B(r)}+\frac{p_{\theta}^{2}}{D(r)}+\frac{J_{z}^{2}}{D(r)\sin^{2}\theta}\Big{)},$ (22)
@@
Geodesic equation for photon $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=0$ is
$-A(r)\dot{t}^{2}+B(r)\dot{r}^{2}+D(r)\dot{\phi}^{2}=0.$ (25)
Now using equation (20) and (25), we get $\dot{r}^{2}/\dot{t}^{2}=(dr/d\phi)$
@@
From the above equation, we can deduced impact parameter $b$ by imposing the condition $dr/d\phi|_{R}=0$.
@@
Now, we define the function
$h^{2}(x)\equiv\frac{D(x)}{A(x)}.$ (28)
@@
When $R$ becomes radius of photon sphere $r_{ph}$ Wang:2022kvg
@@
The critical impact parameter $r_{sh}\equiv b_{cr}$ is equal to $r_{ph}/\sqrt{A(r_{ph})}$. This represented in the figure (1).
@@
This equation is useful to calculate radius of the photon sphere in GR and $f(R)$ gravity.
@@ III.1 Shadow in Schwarzschild space-time for static observer
Here we have the parameter $A(r),B(r),D(r)$ in the Schwarzschild metric (1) as Perlick2021CalculatingBH
@@
In the figure (2), we can distinguished between event horizon ($2m$), photon sphere ($3m$) and black hole shadow ($3\sqrt{3}$) in Schwarzschild spacetime. Angular radius of the black hole shadow for large distance $r_{o}>>m$ is

### v3

@@ Abstract
Black hole shadows are a widespread topic in astrophysics. This paper searches for an optical view of the black hole and the relationship between black hole shadow and photon sphere with curvature. We were inspired by the observations of Sagittarius $A^{*}$ and supermassive black hole M87* through the Event Horizon Telescope. In this paper, we have found the signature of the modified theory of gravity on the photon sphere and shadow. We considered $f(R)$ modified theory of gravity. Our finding is that the new scalar degree of freedom $F$ appears in the expressions of photon sphere and black hole shadow. Both are changed with $F$, and indeed, they depends on spacetime metric.
@@
The first image of the shadow was captured in 2019 by Event Horizon Telescope (EHT) Akiyama et al. (2019a). This is a groundbreaking observation black hole and used for studying the nature of black hole. The observational result of EHT is useful for testing the theory of gravity.
@@
Till now, we do not have direct observational evidences about the dark matter particle. Another mysterious component of the universe is dark energy, which responsible for the accelerated expansion of the universe. There are many methods and theories to solve both mysteries, dark matter and dark energy. The simplest way to tackle these problems by introducing a $\Lambda$ term for late time expansion Riess et al. (1998); Carroll (2001); Sahni (2002); Sahni and Starobinsky (2000); Sahni and Sen (2017) and Cold Dark Matter (CDM) as a candidate of dark matter Bertone et al. (2005); Armendariz-Picon and Neelakanta (2014); Ilić et al. (2021); Perivolaropoulos and Skara (2022); Yang et al. (2025). There are many other way to explain them and one of them is the modification in the curvature. We used $f(R)$ models for accelerated expansion and large scale structure formation Sharma and Verma (2022a, b) and dark matter Yadav and Verma (2019); Sharma et al. (2020a, 2021).
This article seeks to unravel the complexities surrounding black holes in the photon sphere Snepppen (2021), the very first layer of these bodies. These are the boundaries of the Black Hole Shadow (BHS).
@@
On the other hand, shadow is like an illusion of four dimensions of space-time and it depends on the many factors like background metric, static or moving observer and its position etc. Recent picture taken by the Event Horizon Telescope (EHT) of the Sagittarius $A^{*}$ (Sgr A*, pronounced ”sadge-ay-star”) Akiyama et al. (2022).
@@
The trajectory of light bends around the object because of high-density. If light bends so strongly that it moves on circular path because of a high density object like a black hole Snepppen (2021). These orbits of light are known as photon spheres. Dark spots in the middle are termed as black hole shadow. In 1966, Synge studied path of the photons for null geodesics Synge (1966) for Schwarzschild and non-rotating black hole.
@@
In the past few years, many researchers have been investigated the black hole shadows. We know that the size of the black hole shadow is 2.4 times larger than the particle horizon Wang et al. (2022a); Perlick and Tsupko (2021). This introduces some concept about the illusion of black hole shadow. It can be explained with the help of location of the observer. That means BHS depends on observer’s position, accretion disc that means matter around the horizon, rotation of black hole, and charged black hole Takahashi (2004); Li et al. (2020); Gralla et al. (2019). Shapes of BHS strongly depends on modified theory of gravity models. Black hole shadow can be strong feature of the BH that constraint model’s parameters of the modified theory of gravity.
I Observational Evidence
[…]
@@ I Observational Evidence
Additional observational data of the supermassive black hole Sgr A, published by the Event Horizon Telescope (EHT) Akiyama et al. (2022), revealed an angular shadow diameter $\theta_{sh}=48.7\pm 0.7$ $\mu as$ and Schwarzschild shadow deviation $0.08^{+0.09}_{-0.09}$ (VLTI), $0.04^{+0.09}_{-0.10}$ (Keck), with the black hole’s mass $4.0^{+1.1}_{-0.6}\times 10^{6} MM_{\odot}$ Akiyama et al. (2019a).
@@ II Theory background
The black hole shadow and photon sphere can be calculated analytically in general theory of relativity or other theories. These depend on the metric tensor and geodesic for photon. Now, we have the Schwarzschild metric for static, non-rotating, and uncharged symmetric sphere Schwarzschild (1916); Narlikar (2002); Schutz (2022)
@@
The Lagrangian written as Weinberg (1972)
@@
Euler-Lagrange equation is Weinberg (1972); Schutz (2022)
@@
Lagrangian have two cyclic coordinates $t$ and $\phi$ therefore we have two constant of motion ${\partial\mathcal{L}}/{\partial t}=0$ and ${\partial\mathcal{L}}/{\partial\phi}=0$. Because Lagrangian have two cyclic coordinates associated with two conserved quantities energy $E$ and angular momentum of photon $J_{z}$, respectively
@@
where $c=1$. These are very useful quantities to study the trajectories of photon.
@@ II.1 Geodesic and photon orbit
Einstein’s general theory of relativity is most favorable theory for studying black hole der Wissenschaften zu Berlin. (1915). Imagine spacetime as a flexible fabric; massive objects like stars and planets create indentations, and black holes represent points where the fabric is stretched into an infinitesimal point-a singularity. We used Schwarzschild metric equation (1) to calculate photon sphere.
Now, we have calculated the mathematical and geometrical expression of the photon sphere. The Photon moves around the massive object like black hole on the null geodesic i.e. $ds^{2}=0$ or $g_{\mu\nu}dx^{\mu}dx^{\nu}=0$.
Null geodesic equation in the equatorial plane is Zeng et al. (2022)
$\Big{(}1-\frac{R_{s}}{r}\Big{)}dt^{2}+\frac{dr^{2}}{\Big{(}1-\frac{R_{s}}{r}\Big{)}}+r^{2}d\phi^{2}=0.$ (7)
If we divide by the affine parameter $d\lambda$ Zeng et al. (2022)
$\Big{(}1-\frac{R_{s}}{r}\Big{)}\dot{t}^{2}+\frac{\dot{r}^{2}}{\Big{(}1-\frac{R_{s}}{r}\Big{)}}+r^{2}\dot{\phi}^{2}=0.$ (8)
@@
Figure 1: The dotted line is known as the angular diameter of black hole shadow $\alpha_{sh}$ or acceptance angle $r_{sh}=3\sqrt{3}m$ and red circle is the photon sphere $r_{ph}=3m$ for static observer. If observer is at the large distance $r_{o}>>m$, then $r_{sh}$ equal to impact factor Wang et al. (2022b); Perlick and Tsupko (2021). This figure shows that the sizes of black hole horizon, photon sphere and black hole shadow, respectively in Schwarzschild spacetime metric and for static observer
@@ II.2 Effective Potential
The effective potential can be used to calculate the radius of the photon sphere by $dV_{eff}/dr=0$ Zeng et al. (2022). Now calculate the effective potential using the null geodesic equation (8) and using the value of $E$ &\& $J_{z}$ from equations (5) and 6, respectively.
$\frac{E^{2}}{\Big{(}1-\frac{R_{s}}{r}\Big{)}}+\frac{\dot{r}^{2}}{\Big{(}1-\frac{R_{s}}{r}\Big{)}}+\frac{J^{2}}{r^{2}}=0,$ (9)
@@ II.3 Photon Sphere
Radius of the photon sphere calculated from null geodesic, i.e., the geodesic for photon in the equatorial plane. The radius of the photon sphere calculated by using effective potential (11) and differentiating with respect to $r$ Wang et al. (2022a); Perlick and Tsupko (2021)
$\frac{dV_{eff}}{dr}=-\frac{J^{2}}{r^{3}}+\frac{3}{2}\frac{J^{2}}{r^{4}}R_{s}$ (12)
@@
$\frac{J^{2}}{r^{3}}+\frac{3}{2}\frac{J^{2}}{r^{4}}R_{s}=0$ (13)
@@ II.4 Shadow of Black hole
Mathematical explanation of the black hole shadow for the first time given by Synge Synge (1966) in 1996. He considered spherical symmetric Schwarzschild metric foe static universe. He takes $\rho=r/2m$ and $\tau=t/2m$, where $G=c=1$. Constant of motion for angular momentum and energy conservation given as $\rho^{2}\dot{\theta}=\alpha$ and $(1-\rho^{-1})\dot{t}=\alpha\beta$, respectively. Path of the photon $\theta=$ const. &\& $\phi=$const. and Small deviation of light ray (photon) is $\cot\alpha=\sqrt{g_{\rho\rho}/g_{\theta\theta}}(d\rho/d\theta)$ and $\sin^{2}\alpha=(\rho-1)/\rho^{3}\beta^{2}$. We have $\beta=4/27$ for $\rho=3/2$.
@@ III General expression of Black Hole Shadow and photon sphere
Most of the part of this section use the review article Wang et al. (2022a); Perlick and Tsupko (2021). We have find out the general solution of the black hole shadow for spherical symmetric case in the static universe. In this case metric is
@@
where, $A(r)$, $B(r)$ and $D(r)$ are constant. If we use A(r)=12m/rA(r)=1-2m/r, $B(r)=(1-2m/r)^{-1}$ and $D(r)=r^{2}$, $m=G_{N}M/c^{2}$. We find constant of motion equation form (4) as
$-p_{t}=E=-A(r)\dot{t},$ $p_{\theta}=J_{z}=D(r)\sin^{2}\theta\dot{\phi}$ (19)
@@
$E=-A(r)\dot{t},$ $J_{z}=D(r)\dot{\phi}.$ (20)
We have effective potential $V_{eff}$ from Hamiltonian Wang et al. (2022b), $\mathcal{H}=(1/2)g^{\mu\nu}p_{\mu}p_{\nu}$
$\mathcal{H}=\frac{1}{2}\Big{(}\frac{p_{t}^{2}}{A(r)}+\frac{p_{r}^{2}}{B(r)}+\frac{p_{\theta}^{2}}{D(r)}+\frac{p_{\phi}^{2}}{D(r)\sin^{2}\theta}\Big{)},,$ (21)
$\mathcal{H}=\frac{1}{2}\Big{(}-\frac{E^{2}}{A(r)}+\frac{p_{r}^{2}}{B(r)}+\frac{p_{\theta}^{2}}{D(r)}+\frac{J_{z}^{2}}{D(r)\sin^{2}\theta}\Big{)}$ (22)
@@
Geodesic equation for photon $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=0$ is Perlick and Tsupko (2021)
$A(r)\dot{t}^{2}+B(r)\dot{r}^{2}+D(r)\dot{\phi}^{2}=0.$ (25)
Now using equation (20) and (25), we get $\dot{r}^{2}/\dot{t}^{2}=(dr/d\phi)$ Perlick and Tsupko (2021)
@@
From the above equation, we can deduced impact parameter $b$ by imposing the condition $dr/d\phi|_{R}=0$ Perlick and Tsupko (2021).
@@
Now, we define the function Perlick and Tsupko (2021)
$h^{2}(x)\equiv\frac{D(x)}{A(x)}$ (28)
@@
When $R$ becomes radius of photon sphere $r_{ph}$ Perlick and Tsupko (2021)
@@
The critical impact parameter $r_{sh}\equiv b_{cr}$ is equal to $r_{ph}/\sqrt{A(r_{ph})}$. This represented in the fig. (1).
@@
This equation is useful to calculate the radius of the photon sphere in GR and $f(R)$ gravity.
@@ III.1 Shadow in Schwarzschild space-time for static observer
Here we have the parameter $A(r),B(r),D(r)$ in the Schwarzschild metric (1) as Perlick and Tsupko (2021)
@@
In the fig. (1), we can distinguished between event horizon ($2m$), photon sphere ($3m$) and black hole shadow ($3\sqrt{3}$) in Schwarzschild spacetime. Angular radius of the black hole shadow for large distance $r_{o}>>m$ is